OpenStudy (anonymous):

Which of the equations below is equivalent to; `18(1/9x + 1/3y) = 15` A. `2x + 6y = 15` B. `4x + 6y = 270` C. `x + 3y = 15` D. `1/9x + 1/3y = 270`

OpenStudy (akashdeepdeb):

\[18(\frac{ 1x }{ 9 } + \frac{ 1y }{ 3 }) = 15\] \[18(\frac{ x }{ 9 } + \frac{3y }{ 9 }) = 15\] \[2(\frac{ x + 3y }{ 1 }) = 15\] \[2(x + 3y) = 15\] \[2x + 6y = 15\] Understood? :) Ask if you did not. :D